what is the length of the latus rectum of the parabola x^2-6x-2y-1=0?

Answer by stanbon(75887)

what is the length of the latus rectum of the parabola
x^2-6x-2y-1=0?
-----------
Complete the square::
2y+1 = x^2-6x-1
-----
2y+1+9 = x^2-6x+9
(x-3)^2 = 2(y+5
-------
Latus Rectum = 4p = 2
==========================
Cheers,
Stan H.